Hi, can someone help me solve these problems, and show me the steps? Thanks.
1. x^3-x^2-20x=0
2. 5x^2+3x=2
3. x(x-5)=36
This problem is an equation and I have tried to figure out the answer but I'm stuck.
4. 4a^4b+32ab^4

Found 2 solutions by stanbon, checkley71:
Answer by stanbon(75887)

1. x^3-x^2-20x=0
x(x^2-x-20)=0
x(x-5)(x+4)=0
x=0 or x=5 or x=-4
----------------------------
2. 5x^2+3x=2
5x^2+3x-2=0
Use the quadratic formula to get:
x=[-3+sqrt(9-4*5*2)]/10 or x==[-3-sqrt(9-4*5*2)]/10
x=[-3+2sqrt10]/10 or x=[-3-2sqrt10]/10
-------------------------------------
3. x(x-5)=36
x^2-5x-36=0
(X-9)(X+4)=0
x=9 or x=-4
-----------------------------------------
This problem is an equation and I have tried to figure out the answer but I'm stuck.
4. 4a^4b+32ab^4
Not an equation as you have posted it.
Cheers,
Stan H.

Answer by checkley71(8403)

X~3-X~2-20X OR X(X~2-X-20) OR X(X-5)(X+4)=0 OR X=0, X=5 & X=-4
5X~2+3X-2=0 OR (5X-2)(X+1)=0 OR 5X=2 OR X=2/5 & X=-1
X(X-5)=36 OR X~2-5X-36=0 OR(X-9)(X+4)=0 OR X=9 & X=-4
4A~4B+32AB~4 OR 4AB(A~3+8B~3)